Please help me solve this equation.
Find my age, my grandmother's age and my sister's age if:
My father is 40 and he is four years older than my mother
My mother is three times more than my age.
And my sister is 75% my age and 15% of my grandmother's age
Answers:
My age = 40-4/3= 12 years old
My sister's age = 75% of 12 = 9 + 15% of grandmother's age

Answer by rothauserc(4718)

let x be your mother's age
40 = x + 4
let y be your age
x = 3y
solve first equation for x
x = 36
substitute for x in second equation
36 = 3y
3y = 36
y = 12
you are 12 years old
your sister is .75 * 12 = 9 years old
let z be your grandmother's age
.15 * z = 9
z = 9 / .15 = 60
your grandmother is 60 years old